40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Thu Sep 16 15:21:21 2021, Paco, DailyProgress, 1418 nm AUX ECDL, Free space AOM 
    Reply  Wed Sep 22 16:44:34 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, Free space AOM 
       Reply  Mon Oct 4 17:44:34 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, Free space AOM 
          Reply  Tue Oct 19 13:52:03 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1418nm ECDL Frequency noise ECDL_diagram.pdfECDL_FNM_13-10-2021_151524.pdf
             Reply  Tue Mar 8 09:32:56 2022, Paco, DailyProgress, 1418 nm AUX ECDL, 1418nm ECDL Frequency noise schematic.pngraw_bn_spectrum.pngphase_timeseries.pngecdl_freqnoise.png
Message ID: 1926     Entry time: Mon Oct 4 17:44:34 2021     In reply to: 1925     Reply to this: 1927
Author: Radhika 
Type: DailyProgress 
Category: 1418 nm AUX ECDL 
Subject: Free space AOM 

[Paco, Radhika]

Last Friday we received a new lens to direct the AOM 1st-order beam from free space into a fiber cable. We mounted the lens and connected a fiber cable into the photodiode, and tried to align the lens and see a jump in the oscilloscope. We were not able to do so and wrapped up for the day.

Today we continued aligning the lens with the fine adjustment on the mount, and eventually saw signal on the scope! Hooray, done with free space. We then prepared for eventually taking a heterodyne beat note measurement and hooked up the appropriate inputs/outputs to the beamsplitters. We added in the 50-50 beamsplitter that takes in the 1st order diffracted beam along with the beam from the delay line as inputs. We passed one of the outputs to the photodiode and had to retweak the freespace-to-fiber lens until we recovered signal on the scope, and we saw the beatnote signal.

Next, while Paco is out of town I will continue to work towards making a frequency noise measurement. We made a roadmap today:

I will demodulate the beat note using a mixer and a 35 MHz LO sourced from the Marconi. The result will be a 2f cosine term, along with a much lower frequency term which encloses the frequency noise information. This will be passed through a low-pass filter to get rid of the first high-frequency term. The remaining time-domain signal will be passed to the SR785 to obtain a spectra of the frequency noise. Calibration will need to be performed to obtain the right units for the spectra, Hz2/Hz (or Hz/rtHz). 

ELOG V3.1.3-